Weather: Progress report for the month of February

Most of the month we spent waiting for a graphics package, NCAR graphics. We recieved and installed the package on February 28 1996. Since then I have run their example runs and examined the graphics from the output. I have then mailed a message off to Kelvin Droegemeir to see when/if I could go down to Oklahoma. Bob Farallel is trying to help. He contacted Kelvin, and will let us know when he can help out some more. We are going to try to use February 14, 1996 as our initial data, since it was a good day to use for lake effect snow.

I am starting to look at their data output routines. In particular, I am interested in generating moisture plots.